OpenStudy (anonymous):

My answer: 5/2 What is the slope of the line 5x + 2y = 10?

OpenStudy (anonymous):

-5/2**

OpenStudy (reemii):

you can read the slope when you write your equation in the form \(y = ax + b\).

OpenStudy (reemii):

it's \(a\)

OpenStudy (anonymous):

-5/2?

OpenStudy (reemii):

yes, the steps are: \[ \phantom{\iff\ }5x+2y=10 \\ \iff 2y = -5x + 10 \\ \iff y = -(5/2)x + 10/2 \] the coefficient in front of \(x\) is the slope.

OpenStudy (anonymous):

thank you